# Bullion: A Column Store for Machine Learning

### v2

@@ 2.3. Wide Table Projection
This efficiency comes at the cost of flexibility, as the encoding format is fixed and lacks extensive customization options. In contrast to traditional data analytics, machine learning tasks exhibit a simple data access pattern, reading all training data within a specific time period in a batch-oriented manner, eliminating the need for complex indexing and filtering mechanisms. Consequently, the weakened flexibility of Bullion is not a critical concern in training tasks, where the benefits of improved efficiency and faster data access outweigh the limitations in customization and extensibility. Bullion prioritizes efficiency over flexibility for machine learning, significantly speeding up parsing the file metadata compared to Parquet and ORC, making it a suitable choice for optimizing data retrieval and processing in machine learning training workflows.
@@ 2.4. Storage Quantization
Feature and data representation constitute fundamental determinants of model performance in both recommender systems and LLMs. In recommender systems, the dimensionality of the feature space directly correlates with the size of embedding tables, which in turn determines the model’s total memory footprint. However, production deployment scenarios frequently impose strict storage constraints that inhibit the incorporation of additional features and expanded embedding capabilities. As a consequence, numerous promising architectural improvements remain unrealized in production environments. The impact of these limitations is particularly acute in high-stakes applications such as advertising platforms and AI search engine, where rich feature representations and embedding directly influence business outcomes.
@@
Model quantization Han et al. (2015); Jacob et al. (2018); Deng et al. (2020), a fundamental technique in model generation and deployment, aims to address model size, inference speed, and memory efficiency challenges. It typically reduces model weights from high-precision 32-bit floating-point (FP32) representations to more compact numerical formats, including lower-precision floating-point (FP16, FP8) or integer representations in Figure 6. The proven effectiveness of model quantization in training and serving stages has motivates its application for features and embeddings in storage, thus adapting model quantization to storage quantization. Our key insight is that different features and embeddings exhibit varying degrees of precision sensitivity, enabling a mixed-precision quantization strategy that can be dynamically tuned at the granularity of individual features. The resulting storage savings can be strategically reinvested to enhance model capabilities through expanded sequence lengths and increased embedding dimensionality, ultimately achieving both infrastructure cost reduction and improved model performance in production environments.
@@
Recommender Systems operates on three distinct feature types, each with specific data representations: 1) Dense features, characterized by continuous FLOAT or DOUBLE numerical values, include both direct measurements (e.g., demographic attributes, engagement metrics) and derived features generated through upstream offline model inference. These dense representations typically manifest as floating-point. 2) Sparse features represent high-cardinality categorical variables, predominantly implemented through efficient lookup mechanisms such as INT or BIGINT foreign key references, where the feature space exhibits significant sparsity (e.g., user interaction histories, entity identifiers). 3) Embedding features emerge as learned dense vector representations, typically encoded as FLOAT32 or FLOAT64 arrays, transformed from either dense or sparse inputs through parameterized embedding layers during model training.
@@
While emerging startups in LLM-based search engines (e.g., Perplexity) can leverage third-party search APIs like Google and Bing for result aggregation and LLM-based summarization, major technology companies face distinct constraints that necessitate building proprietary search infrastructure. These constraints include limitations on query volumes, restrictions on logging search results for model training, and stringent latency requirements. Consequently, these companies must implement sophisticated crawling systems that operate at varying frequencies (minute/hour/day) based on domain priorities, while managing substantial computational resources for preprocessing and generating embeddings using offline ML models for similarity-based retrieval. Considering that even Google’s index encompasses hundreds of trillions of pages, the storage of embeddings and associated metadata (such as entity attributes) poses significant capacity challenges, making vector embedding quantization techniques crucial for columnar storage systems.
@@
For integer features, quantization provides lossless compression by rehashing the input space to a smaller range (e.g., INT8, INT16, INT32). For low cardinality columns, column stores can further leverage bit-packed encoding and RLE to achieve higher compression ratios. Native support for low-precision data types not only saves storage space, but also reduces network bandwidth and preprocessing computation overhead. This feature is complementary to encodings like RLE, working orthogonally. Given that sparse features (integers) often constitute a significant portion of the total storage footprint in recommender systems, new columnar formats should incorporate nested light-weight adaptive encoding schemes Kuschewski et al. (2023); nim (2024); Chattopadhyay et al. (2019) to improve compression while minimizing decoding overhead. For float-point features and embeddings, recent insights reveal that the precision offered by 4-byte floats is often superfluous for effective results. Feature quantization, therefore, involves storing values in FP16, BF16, or FP8 formats, usable directly in training and serving, or upcasted to FP32 as an interim measure pending native computational support. This reduction to 1 or 2 bytes per float can halve or quarter storage costs, disk I/O, network bandwidth, decompression efforts, and potentially computing needs during training and inference.
@@
Challenges. First, while native support for the reduced-precision formats (FP16, BF16 and BF8) in data processing frameworks (e.g., Apache Spark, Arrow, Velox) would enhance computational and memory efficiency during feature preprocessing, current implementations require interim solutions where data ingestion pipelines must implement automatic padding mechanisms to convert BF16 to FP32 format. Second, given that embedding vectors are typically normalized to (-1, 1), there remains a need for more storage-efficient encoding schemes specifically optimized for BF16’s numerical characteristics. Third, the primary technical barrier stems from potential degradation in model performance metrics for business-critical applications, particularly in floating-point quantization; one proposed mitigation strategy involves decomposing FP32 values into dual FP16 representations stored across separate columns, with subsequent 1:1 join operations during data retrieval.
@@ 2.5. Multimodal Storage
As introduced in Section 1, LLM pre-training increasingly demands multi-modal data integration, encompassing text, images, audio, and video content. These modalities typically appear in contextually related forms, such as within web pages, serving as collective inputs for model pre-training. However, the conventional column store face challenges in supporting multi-modal data, particularly due to the substantial storage requirements of high-resolution video content.
@@
Challenges. The solution implements a dual-table architecture where meta tables leverage columnar format for efficient metadata management (including text, image and even audio), while media tables adopt row-oriented storage for multi-modal content (specifically video data). While this hybrid storage approach enables comprehensive data organization, it introduces a critical performance bottleneck during training. The separation of data across different storage formats (columnar for metadata and row-oriented for media) and storage media results in fragmented I/O operations, significantly impacting training I/O throughput and overall system performance.
@@
Solution. While specialized diffusion transformer like Sora sor (2024) require high-resolution, full-size videos for generating complex scenes with multiple characters, specific motion patterns, and accurate subject-background details, the pre-training of general-purpose LLM adopts a more efficient approach as shown in Figure 7. The training process can effectively operate on a strategic subset of video frames at reduced resolution, with these critical frames directly integrated into the columnar format. This optimization enables the training system to access text, audio, and video modalities through the columnar storage alone, satisfying the majority of training scenarios. Such an integrated storage strategy eliminates the latency overhead associated with external, fragmented I/O operations. When full-size video access is required, the system still maintains the capability to perform external lookups through video indices stored in the meta table, providing a complete yet flexible data access solution.
@@
The meta table faces an additional challenge of random I/O patterns during training, as only high-quality data samples are selected for the training process. The quality-based filtering mechanism, primarily driven by quality scores, typically excludes a significant portion of the data, resulting in non-sequential access patterns within the columnar storage. To address this I/O inefficiency, the storage system implements a quality-aware data organization strategy: incoming row data is presorted by quality score in descending order prior to insertion into the storage. This presorting approach ensures contiguous access to high-quality video frames during training, effectively minimizing random I/O operations and optimizing storage access patterns.
@@
The quality-based reordering strategy for LLM training represents a row reordering approach, where entire rows are sorted based on their quality scores, ensuring that high-quality video frames and their associated features are stored contiguously within pages of each row group. This contrasts with the column-oriented reordering requirements in recommendation systems, where typically only 10% of the thousands of available features are accessed during training. In recommendation scenarios, the system prioritizes frequently accessed, important features through column reordering, ensuring these features (columns) are stored contiguously within row groups. While the LLM training optimization focuses on row-level sorting based on quality scores, leading to sequential access of high-quality multi-modal data, the recommendation system optimization emphasizes column-level importance sorting. Both approaches achieve their respective performance gains by minimizing random I/O through strategic data organization, though they operate on different dimensional axes of the storage structure.
Traditional data organization typically relies on query engines in data lakes or lakehouses to perform row-wise or column-wise sorting of existing columnar data, whereas machine learning workloads demand more specialized capabilities. The columnar storage format itself should provide native interfaces for data organization during the write path, enabling both row-wise and column-wise runtime sorting through user-defined functions (UDFs) at write time.
@@ 2.6. Cascading Encoding
Unlike the previously discussed challenges, cascading encoding represents a general-purpose solution applicable to all columnar storage scenarios. Given that machine learning workloads predominantly consist of integer and floating-point data types, the runtime adaptive selection of nested encoding schemes for each column proves highly effective. Recent work, including BtrBlocks Kuschewski et al. (2023) and Nimble nim (2024), has adopted cascading encoding approaches but exhibits two limitations. First, existing formats each implement only a subset of available encoding schemes, with no single implementation providing comprehensive coverage, thus failing to achieve optimal performance across different scenarios. Second, current Parquet and ORC tightly couple various encoding methods (such as RLE, varint, and fixed-width encoding) without providing unified interfaces, making it impossible to utilize these encoding schemes independently. This coupling presents a significant barrier to implementing cascading encoding, which requires modular encoding selection–a key motivation driving the community’s discussions on Parquet v3 evolution par (2024). Based on these observations, Bullion proposes a comprehensive encoding catalog in Table 2 and advocates for an independent encoding module–towards functional decomposition–that provides cascading encoding capabilities. Such a modular approach would enable all columnar storage formats to leverage diverse encoding methods effectively.
Table 2. Comprehensive overview of column encoding scheme. This comprehensive encoding catalog is derived from analyzing various storage systems and formats, including Parquet, ORC, Kudu, BtrBlocks Kuschewski et al. (2023), Nimble nim (2024), Protocol Buffers, and other columnar storage implementations.
@@
The composable and recursive nature of encodings enables powerful combinatorial patterns that can achieve superior data compression compared to traditional static, single-encoding approaches. As the encoding catalog expands, existing encodings become more effective by gaining access to a broader selection of subcomponent options. However, this flexibility introduces new challenges. First, the search space for optimal encoding combinations grows significantly, requiring systems like Procella Chattopadhyay et al. (2019) and BtrBlocks Kuschewski et al. (2023) to employ sampling-based distribution analysis and heuristic approaches for encoding selection. Second, the unbounded recursive potential raises an important open question regarding the optimal depth of recursion. Current implementations, such as BtrBlocks, pragmatically limit recursion to one or two levels, but determining the ideal recursion depth remains an active research topic.
While prior work Zeng et al. (2023) suggests that "formats should not apply general-purpose block compression by default", we present a contrasting perspective based on our empirical observations. As evidenced by the Chunked encoding in Table 2, which leverages zstd for byte-level compression, general-purpose compression methods continue to demonstrate practical value in specific scenarios. This is particularly evident in time-series data where many columns are rarely accessed, and in recommendation systems where only approximately 10% of thousands of features are frequently queried—scenarios where block compression proves highly effective for storage optimization.
@@ 3. Related Work
Meta has unveiled Alpha Chattopadhyay et al. (2023); Zhao et al. (2022), a novel columnar format aimed at boosting wide-table projection efficiency, especially for ML use cases, addressing significant weaknesses in previous formats that include: 1) feature flattening, which stores each feature as a separate stream on the disk—effectively treating n features as n columns instead of a single map. By using a ’flat map’ data type, where keys and values are in distinct columns, this approach significantly aids ML training processes that often need just a subset of features, enhancing wide-table projection efficiency. 2) feature reordering, which arranges commonly accessed features in adjacent disk positions to minimize over-reads from storage. 3) Coalesced reads, which bundle selected feature streams into single I/O operations of up to 1.25 MiB, thus eliminating storage throughput degradation by amortizing disk seeks. Despite these advancements, Alpha does not explicitly address the metadata overhead challenge, data compliance, sparse feature encoding, and feature quantization, which are primary focuses of Bullion.
Recent advances in columnar storage have emerged from both academia and industry, with notable contributions from BtrBlocks Kuschewski et al. (2023) and Nimble nim (2024), which draw inspiration from Google’s Procella Chattopadhyay et al. (2019). These new formats revisit the concept of cascading/nested encodings at runtime, albeit with different optimization objectives. BtrBlocks primarily focuses on optimizing decoding efficiency through its nested encoding scheme, while Nimble introduces a more comprehensive and granular approach to encoding selection. Nimble’s innovation lies in its user-configurable linear objective function that independently weights read time, write time, and storage size. This flexible optimization framework enables users to tailor encoding strategies to their specific workload requirements, whether prioritizing query performance, compression efficiency, or write throughput. By evaluating each potential encoding against these weighted metrics, Nimble can automatically select the optimal encoding strategy that minimizes the overall cost according to the user-defined objective function, representing a significant advancement in adaptive columnar storage systems.
While these studies provide valuable insights, they do not comprehensively address the storage-level challenges and opportunities unique to machine learning workloads. Bullion distinguishes itself by directly targeting these emerging challenges and opportunities, aiming to bridge the gap between columnar storage and the evolving demands of machine learning applications.

### v3

@@ 2.3. Wide Table Projection
This efficiency comes at the cost of flexibility, since the encoding format is fixed and lacks some customization options. This lack of flexibility is less problematic for machine learning tasks. since they often exhibit a simple data access pattern, reading all training data within a specific time period in a batch-oriented manner, without requiring complex indexing or filtering. Thus, the benefits of improved efficiency and faster data access outweigh the limitations in customization and extensibility.
@@ 2.4. Storage Quantization
Feature and data representation are important determinants of model performance in both recommender systems and LLMs. In recommender systems, the dimensionality of the feature space directly correlates with the size of embedding tables, which in turn determines the model’s total memory footprint. However, production deployment scenarios frequently impose strict storage constraints that inhibit the incorporation of additional features and expanded embedding capabilities. As a consequence, numerous promising architectural improvements remain unrealized in production environments. This is particularly problematic for advertising platforms and AI search engines, where rich feature representations and embedding directly influence business outcomes.
@@
Model quantization Han et al. (2015); Jacob et al. (2018); Deng et al. (2020) aims to address model size, inference speed, and memory efficiency challenges. It typically reduces model weights from high-precision 32-bit floating-point (FP32) representations to more compact numerical formats, including lower-precision floating-point (FP16, FP8) or integer representations (see Figure 6). The proven effectiveness of model quantization in training and serving stages motivates its application for features and embeddings in storage, adapting model quantization to storage quantization. Different features and embeddings exhibit varying degrees of precision sensitivity, which implies that a mixed-precision quantization strategy should be used that can be dynamically tuned at the granularity of individual features. The resulting storage savings can be strategically reinvested to enhance model capabilities through expanded sequence lengths and increased embedding dimensionality, ultimately achieving both infrastructure cost reduction and improved model performance in production environments.
@@
Recommender Systems operate on three distinct feature types, each with specific data representations: 1) Dense features are characterized by continuous FLOAT or DOUBLE numerical values, include both direct measurements (e.g., demographic attributes, engagement metrics) and derived features generated through upstream offline model inference. 2) Sparse features represent high-cardinality categorical variables, predominantly implemented through efficient lookup mechanisms such as INT or BIGINT foreign key references, where the feature space exhibits significant sparsity (e.g., user interaction histories, entity identifiers). 3) Embedding features are learned dense vector representations, typically encoded as FLOAT32 or FLOAT64 arrays, transformed from either dense or sparse inputs through parameterized embedding layers during model training.
@@
Many existing technology companies face constraints from third-party search APIs (e.g., from Google or Bing) that motivate them to build their own proprietary search infrastructure for result aggregation and LLM-based summarization. These constraints include limitations on query volumes, restrictions on logging search results for model training, and stringent latency requirements. Consequently, they implement sophisticated crawling systems that operate at varying frequencies (minute/hour/day) based on domain priorities, while managing substantial computational resources for preprocessing and generating embeddings using offline ML models for similarity-based retrieval. Since the crawled data may include hundreds of trillions of pages, the storage of embeddings and associated metadata (such as entity attributes) poses significant capacity challenges, thereby increasing the motivation for vector embedding quantization techniques.
@@
For integer features, quantization provides lossless compression by rehashing the input space to a smaller range (e.g., INT8, INT16, INT32). For low cardinality columns, column stores can further leverage bit-packed encoding and RLE to achieve higher compression ratios. Native support for low-precision data types not only saves storage space, but also reduces network bandwidth and preprocessing computation overhead. Given that sparse features (integers) often constitute a significant portion of the total storage footprint in recommender systems, new columnar formats should incorporate nested light-weight adaptive encoding schemes Kuschewski et al. (2023); nim (2024); Chattopadhyay et al. (2019) to improve compression while minimizing decoding overhead. For float-point features and embeddings, the precision offered by 4-byte floats is often unnecessary for effective results. Feature quantization therefore stores values in FP16, BF16, or FP8 formats, usable directly in training and serving. This reduction to 1 or 2 bytes per float can halve or quarter storage costs, disk I/O, network bandwidth, decompression efforts, and compute cost during training and inference.
@@
Opportunities. There are three potential areas that could improve the performance of the approach discussed above. First, native support for the reduced-precision formats (FP16, BF16 and BF8) in data processing frameworks (e.g., Apache Spark, Arrow, Velox) would enhance computational and memory efficiency during feature preprocessing (current implementations require interim solutions where data ingestion pipelines must implement automatic padding mechanisms to convert BF16 to FP32 format). Second, given that embedding vectors are typically normalized to (-1, 1), there exists a need for more storage-efficient encoding schemes specifically optimized for BF16’s numerical characteristics. Third, some FP32 features are crucial for business-critical models. To mitigate potential accuracy degradation from FP16 quantization while maintaining computational efficiency, it is possible to use a dual-column storage strategy: decomposing FP32 features into two FP16 representations. This approach enables business-critical models to reconstruct original FP32 precision through 1:1 join operations during feature retrieval, while allowing other models to utilize FP16 features. Although this strategy does not reduce storage footprint, it decreases network bandwidth utilization and computational overhead in upstream processing stages.
@@ 2.5. Multimodal Storage
As stated in Section 1, LLM pre-training increasingly demands multi-modal data integration, encompassing text, images, audio, and video content. These data types typically appear in contextually related forms, such as within web pages, serving as collective inputs for model pre-training. Meanwhile, conventional column stores face challenges in supporting multi-modal data, especially for high-resolution video content.
@@
Challenge. As described in Section 1, it is possible to use a dual-table architecture where meta tables leverage columnar format for efficient metadata management (including text, image and even audio), while media tables adopt row-oriented storage for multi-modal content. However, this approach introduces a performance bottleneck during training. The separation of data across different storage formats (columnar for metadata and row-oriented for media) and locations results in fragmented I/O operations, significantly impacting training I/O throughput and overall system performance.
@@
Solution. While specialized diffusion transformers like Sora sor (2024) require high-resolution, full-size videos for generating complex scenes with multiple characters, specific motion patterns, and accurate subject-background details, the pre-training of general-purpose LLMs does not require this same level of resolution. Instead, it can effectively operate on a strategic subset of video frames at reduced resolution, with these critical frames directly integrated into the columnar format, as shown in Figure 7. This optimization enables the training system to access text, audio, and video modalities through the columnar storage alone, satisfying the majority of training scenarios. By having all the required data stored together, this approach eliminates the latency overhead associated with external, fragmented I/O operations. When full-size video access is required, the system still maintains the capability to perform external lookups through video indices stored in the meta table.
@@
The meta table faces an additional challenge of random I/O patterns during training, as only high-quality data samples are selected for the training process. The quality-based filtering mechanism, primarily driven by quality scores, typically excludes a significant portion of the data, resulting in non-sequential access patterns within the columnar storage. To alleviate this I/O inefficiency, the storage system implements a quality-aware data organization strategy: incoming row data is presorted by quality score in descending order prior to insertion into the storage. This presorting approach improves contiguous access to high-quality video frames during training.
@@
The quality-based reordering strategy for LLM training is a row reordering approach, where entire rows are sorted based on their quality scores, ensuring that high-quality video frames and their associated features are stored contiguously within pages of each row group. This contrasts with the column-oriented reordering requirements in recommendation systems, where typically only 10% of the thousands of available features are accessed during training. In recommendation scenarios, the system prioritizes frequently accessed, important features through column reordering, ensuring these features (columns) are stored contiguously within row groups. Both approaches achieve their respective performance gains by minimizing random I/O through strategic data organization, though they operate on different dimensional axes of the storage structure.
@@ 2.6. Cascading Encoding
Given that machine learning workloads predominantly consist of integer and floating-point data types, adaptively selecting nested encoding schemes for each column at runtime can be highly effective. Recent work, including BtrBlocks Kuschewski et al. (2023) and Nimble nim (2024), have adopted cascading encoding approaches; however, existing formats implement only a subset of available encoding schemes (listed in Table 2), with no single implementation providing comprehensive coverage. Furthermore, Parquet and ORC currently tightly couple various encoding methods (such as RLE, varint, and fixed-width encoding) without providing unified interfaces, making it impossible to utilize these encoding schemes independently. This coupling presents a significant barrier to implementing cascading encoding, which requires modular encoding selection. What is needed is an independent encoding module that provides cascading encoding capabilities. Such a modular approach would enable all columnar storage formats to leverage diverse encoding methods effectively.
Table 2. Catalog of column encoding schemes found in existing storage systems and formats including Parquet, ORC, Kudu, BtrBlocks Kuschewski et al. (2023), Nimble nim (2024), and Protocol Buffers.
@@
The composable and recursive nature of encodings enables combinatorial patterns that can achieve superior data compression compared to static, single-encoding approaches. However, the search space for optimal encoding combinations grows significantly as the catalog expands, requiring systems like Procella Chattopadhyay et al. (2019) and BtrBlocks Kuschewski et al. (2023) to employ sampling-based distribution analysis and heuristic approaches for encoding selection. The unbounded recursive potential raises an important question regarding the optimal depth of recursion. Current implementations, such as BtrBlocks, pragmatically limit recursion to one or two levels, but determining the ideal recursion depth still requires investigation.
While prior work Zeng et al. (2023) suggests that "formats should not apply general-purpose block compression by default", general-purpose compression methods (such as Chunked encoding in Table 2) continue to demonstrate practical value in specific scenarios. This is particularly evident in time-series data where many columns are rarely accessed, and in recommendation systems where only approximately 10% of thousands of features are frequently queried—scenarios where block compression proves highly effective for storage optimization.
@@ 3. Related Work
Meta has unveiled Alpha Chattopadhyay et al. (2023); Zhao et al. (2022), a novel columnar format aimed at boosting wide-table projection efficiency, especially for ML use cases, addressing significant weaknesses in previous formats. It includes: 1) feature flattening, which stores each feature as a separate stream on disk—effectively treating n features as n columns instead of a single map. 2) feature reordering, which arranges commonly accessed features in adjacent disk positions to minimize over-reads from storage. 3) coalesced reads, which bundle selected feature streams into single I/O operations of up to 1.25 MiB, thereby amortizing disk seeks and improving throughput. Despite these advancements, Alpha does not explicitly address the metadata overhead challenge, data compliance, sparse feature encoding, and feature quantization, which are primary focuses of Bullion.
Recent advances in columnar storage have emerged from both academia and industry, including BtrBlocks Kuschewski et al. (2023) and Nimble nim (2024), which draw inspiration from Google’s Procella Chattopadhyay et al. (2019). These new formats revisit the concept of cascading/nested encodings at runtime, albeit with different optimization objectives. BtrBlocks primarily focuses on optimizing decoding efficiency through its nested encoding scheme, while Nimble introduces a more comprehensive and granular approach to encoding selection. Nimble incorporates a user-configurable linear objective function that independently weights read time, write time, and storage size that enables users to tailor encoding strategies to their specific workload requirements. These research efforts focus on improving data encoding, but do not comprehensively address the storage-level challenges and opportunities unique to machine learning workloads which is the focus of Bullion.